LOW TEMPERATURE HELIUM PRESSURE BROADENING OF AMMONIA INVERSION TRANSITIONS, 10 - 35K

\underline{DANIEL R. WILLEY}, P. DENARDO, P. NULL, P. PONDILLO AND T. TYSZKA, Department of Physics, Allegheny College, Meadville, PA 16335.

We report low temperature helium pressure broadening cross sections of NH_{3} inversion transitions from a collisionally cooled experiment. Helium pressure broadening data was obtained for the three microwave inversion transitions of the (J, K) rotational levels, (1, 1), (2, 2) and (3, 3), over a temperature range of 10 - 35 K. Pressure broadening cross sections were also calculated for comparison with the experimental data using three existing NH_{3}-He potential surfaces. The agreement between experiment and theory is not particularily good, especially for the ortho-NH_{3} level, (3, 3), and appears to point out the need for further work in either the description of the potential surface or the modeling of the collision dynamics.